# Timing characterization of MALTA and MALTA2 pixel detectors using Micro X-ray source

### v1

@@ Abstract
The MALTA monolithic active pixel detector is developed to address the challenges anticipated in future high-energy physics detectors. As part of its characterization, we conducted fast-timing studies necessary to provide a figure of merit for this family of monolithic pixel detectors. MALTA has a metal layer in front-end electronics, and the conventional laser technique is not suitable for fast timing studies due to the reflection of the laser from the metallic surface. X-rays have been employed as a more effective alternative for penetration through these layers. The triggered X-ray set-up is designed to study timing measurements of monolithic detectors. The timing response of the X-ray set-up is characterized using an LGAD. The timing response of the MALTA and MALTA2 pixel detectors is studied, and the best response time of MALTA2 pixel detectors is measured at about 2 ns.
@@ 1 Introduction
Pixel detectors are crucial in particle physics experiments due to their ability to precisely capture and record the individual locations of particle interactions[1]. Their high spatial resolution and fine granularity contribute to improved tracking, help in analyzing complex event topologies, and enhance the overall performance of experiments in unraveling fundamental aspects of particle behavior and interactions. The increasing energy and luminosity of the Large Hadron Collider in CERN[2][3] requires radiation hard pixel detectors[4] to sustain the environment. Therefore, there is a constant effort to make efficient and thinner pixel detectors that would be suitable for high luminosity collisions in terms of radiation hardness, high detection efficiency with low noise, dead time, and delay time. Monolithic silicon pixel detectors [5][6] are found to have better precision, better radiation hardness, and lower noise than traditional bump-based detectors with large area production, lower cost, and lower material usage. Such an attempt led to the development of the MALTA[7] [8]and its second generation MALTA2[9][10] pixel detectors aiming for HL-LHC. All the characteristics of the detector are to be studied before its use in an actual detector. One such property is fast-timing, which is the delay or time taken between the collision of a particle with the detector and its detection. In this study, a Micro-X-ray source is designed and used to study the Fast timing properties of MALTA and MALTA2 detectors.
2 MALTA and MALTA2 Pixel detectors
[…]
@@ 2 MALTA and MALTA2 Pixel detectors
MALTA2,[12] [13]a pixel detector demonstrator, is designed using Tower-Jazz 180 nm Monolithic CMOS technology, measuring $20.2\,\text{mm}\times 10.1168\,\text{mm}$. It is approximately half the size of MALTA, featuring a matrix of $224\times 512$ pixels, with each pixel measuring $36.4\,\mu\text{m}^{2}$. Two process modifications(NGAP and extra deep PWELL/XDPW) has been done to increase the in the pixel corners[14]. It is specifically designed to increase the radiation hardness [15] with achieving lower RTs noise in the sensor front-end. There are improvements in MALTA2 slow control by using a shift resistor instead of an Ethernet-like protocol. In the pixel-front end, the implementation of an open loop amplification led to a compact design and low noise and the use of multiple cascode transistors helped in the gain enhancement of the front end[9]. The demonstrator’s asynchronous read-out is clock-independent, transmitted through a $37$-bit bus, and controlled via a $4322$-bit shift register. The read-out includes KC705, a KINTEX7 series of FPGA[16].
@@ 3 Micro-X-ray set-up at CERN
With the bunch-crossing clock of the LHC being 25 ns[17], sensors and read-out architecture have to be sufficiently fast in order to match the hits with the corresponding bunch-crossing. Furthermore, nanosecond level response time resolution is required to contribute to event reconstruction. The fast timing of the chip requires examination prior to their usage in a detector. Traditionally, this investigation involves reflecting a laser beam off the detector and analyzing the discrepancy between the input and output signals[18]. However, this laser technique proves inadequate for the MALTA2 due to significant reflection issues with the detector’s front-end electronics[7], which interfere with accurate timing measurements.
Figure 1: Figure illustrates the devices used for the study and the connections between them. In (a), The upper left plot shows the complete set-up with the representation of a laser, a High voltage source for X-ray generation, and an oscilloscope. In the left plot in (b), the Kintex7 KC705 FPGA is shown, which is used to trigger the laser and also helps us in signal processing. The right figure in (b) shows the MALTA sensor with the designed micro X-ray tube, held by a linear stage 10 cm above the chip.
To circumvent these challenges, we have opted for X-rays because of their ability to penetrate the metal layers of the detector without reflection. This does not require a high-end X-ray machine. Instead, we have developed a micro X-ray system, as shown in Figure 1, capable of emitting X-rays in response to laser exposure. This approach employs a micro X-ray source that generates short pulses from Cu-Cr targets specifically designed for this application. For the purpose of characterizing the timing performance of the MALTA2, our triggered X-ray set-up plays an important role. Initially, the timing response of the X-ray itself is measured using the LGAD [19]. This baseline measurement serves as a reference for further comparison with the MALTA2’s timing response, enabling us to achieve a comprehensive understanding of the detector’s timing capabilities.
@@ 3.1 Micro-X-ray source
X-rays are generated by utilizing a $16$ mm elongated tube featuring a photo-cathode and anode composed of chromium and copper fabricated above a Beryllium window. Electrons are liberated from the photo-cathode via the photoelectric effect initiated by a laser beam. Subsequently, these emitted electrons undergo acceleration through the application of a $12kV$ electric field within the tube. The interaction of these accelerated electrons with the target material in the anode produces X-rays, including bremsstrahlung and fluorescence. The X-ray production process is illustrated in Figure 2. The emitted X-rays exhibit a random trajectory in space. With the increased applied voltage to the X-ray tube, the number of generated photons increases due to the formation of new transition lines.
Figure 2: Illustrative description of Micro X-ray source. It is connected to a 530 nm laser source that falls on the cathode. The anode material is made with 250 nm copper and 50 nm chromium, which generates x-rays after interacting with accelerated electrons. Finally, there is a beryllium window that prevents electrons from passing through it.
@@ 3.2 Geant4 Simulation for X-ray tube
Figure 3: Geant4 Simulation of the photon energy generated from the Micro X-ray tube. The x-axis shows the energy of the generated photons in the keV scale. Four peaks are visible, with two more prominent peaks at 5.41 and 8.04 keV corresponding to $k_{\alpha}$ energy peaks of Cr and Cu.
A Geant 4 simulation was conducted to analyze the energy variation of emitted photons from a X-ray tube. The simulation results in Figure 3 show distinct peaks at 5.4 and 8 keV, corresponding to the characteristic lines ($k_{\alpha}$) of copper and chromium along with two distinct $k_{\beta}$ lines. Subsequently, experimental validation of these findings was undertaken using an LGAD (Low Gain Avalanche Detector) because of its high granularity and precise timing capabilities.
@@ 4 Test set-up with LGAD
LGAD, the Low-Gain Avalanche Detector[19][20] is a type of silicon detector used in particle physics experiments. The utilization of LGAD in this experimental phase is particularly advantageous, as it enables us to achieve optimal timing resolutions[21], a crucial factor in accurately characterizing X-ray emission. The LGAD experimental arrangement is illustrated in the accompanying schematic diagram, as shown in Figure 4. In this set-up, laser beams are generated and guided onto a micro X-ray source, initiating the generation of X-rays. The resultant X-rays are directed toward the LGAD, and the corresponding signals are captured and recorded using an oscilloscope. Simultaneously, we have directly utilized the LGAD signal to access the energy of the emitted photons based on the amplitude of the LGAD signal. This approach provides a direct and efficient mean of evaluating the photon energy within our experimental framework.
@@
Figure 4: LGAD set-up illustrating the working flow of the set-up. In (a), we are focusing the laser directly on LGAD, and in (b), an additional Micro-Xray source is used. The laser emission and detection time information are collected and compared in both cases using the oscilloscope.
@@ 5 Results from LGAD
We conducted two key studies utilizing the Low Gain Avalanche Diode. The first study focused on characterizing the energy peaks generated by the X-ray source. This aspect is necessary to understand the X-ray tube’s performance and ensure its suitability for our applications. The second study delved into analyzing the timing precision of the X-ray source by examining the standard deviation of its delay. For the same, we compared the performance of the micro X-ray source, activated by laser exposure, against a standalone laser source. These preliminary analyses are important in evaluating the contribution of the X-ray set-up toward total delay to ensure precise timing measurements in our experimental framework.
@@ 5.1 Characterization of Micro X-ray tube
Figure 5: Comparison of the signal amplitudes using LGAD with $Fe^{-55}$ and X-ray sources as a function of voltage measured from the LGAD. The two peaks are visible for the signal using X-rays generated at 12kV supply voltage, corresponding to Cr and Cu $K_{\alpha}$ lines. The peaks of the red line representing the amplitude of the signal using an X-ray tube at 12kV are compared with the peak from the $Fe^{-55}$-source(blue)
In this study, we quantify the energy of photons using the LGAD signal, represented on the X-axis in terms of voltage and exhibit a linear relationship with energy. We employed a $Fe^{-55}$ source with known energy associated with alpha decay to establish a calibration framework. The measured energy from the alpha decay was subsequently compared with the energies of X-rays generated at voltages of 5.5 kV, 8.5 kV, and 12 kV. Notably, the X-rays produced under a 12 kV voltage setting revealed distinct peaks at 5.4 kV and 8 kV, explaining the nuanced energy characteristics inherent in the X-ray emission process. With the increasing energy of the X-ray photon, the energy of the electron emitted increases, increasing the signal voltage. A comparison of peaks is shown in Figure 5. The peak is parametrized using Gaussian and polynomial functions, as shown in Figure 6. The results from the fit are summarized in Table 1, which matches the theoretical value of $k_{\alpha}$ lines from chromium and copper.
Figure 6: The subplots show a comparison between the fitting parameters of generated signal amplitude using LGAD with Fe and X-ray sources at 5.5 keV, 8.5 keV, and 12 keV. The x-axis represents the energy of generated electrons from LGAD using the sources, while the y-axis represents the number of events. The two peaks are visible for the signal amplitude using X-rays generated at 12 kV supply voltage.
@@
Peak Value Energy (Theoretical value) Energy (Observed value)
@@
Peak from 8.5 keV Cr 0.0188 5.41 keV 5.54 keV
Peak from 8.5 keV Cr 0.0270 8.04 keV 7.96 keV
@@ 5.2 Timing performance of LGAD
In this investigation, the LGAD was characterized by utilizing a laser beam, yielding a timing resolutions below 50ps, where the resolution has contributions from both LGAD and the laser source. Subsequently, the entire system was engaged to generate X-rays through laser irradiation. Through analysis of the X-ray signals, the timing resolution is determined to be approximately 300ps across varying laser frequencies, which can be seen in Figure 7. This evaluation allowed us to draw a conclusive inference: the timing resolution of the Micro-X-ray source is estimated to be in the proximity of 250ps, demonstrating its efficiency and reliability in temporal measurements. These findings underscore the potential of our system for applications demanding precise timing resolution in the realm of X-ray generation and detection.
Figure 7: Comparison of the timing delay between the laser trigger and LGAD signal detection time using laser and X-rays. The laser frequency is varied from 5 to 15 MHz, and the RMS of the delay are compared with and without using a micro-X ray source. The RMS of the timing delay using the micro-XRay tube with laser is found to be approximately 250 ps higher than the RMS of timing delay without using the micro-XRay tube, concluding the contribution of approximately 250 ps from the micro-XRay tube towards the RMS of the delay.
@@ 6.1 Measurements using MALTA sample
In contrast to the LGAD set-up, the MALTA system employs an internal trigger from the FPGA to initiate the laser, which is subsequently directed toward the X-ray tube for X-ray emission as described in Figure 8. The operational sequence involves triggering the FPGA using software, whereby the KINTEX-7 FPGA activates the laser source. As the laser beam interacts with the X-ray tube, X-rays are emitted and directed toward the MALTA sensor (W11R12, which is a 100 nm thick sensor with a p-type Czochralski substrate with an extra deep p-well implant). The timing information is gathered within the FPGA and transmitted to the oscilloscope. The oscilloscope then performs a comparison of the timing data, enabling the determination of delay information. This delay information is necessary for analyzing the temporal characteristics of MALTA where we can subtract the effect of previously measured micro X-ray source, which is approximately 300 ps. Notably, the timing performance of MALTA is observed to operate on a nanosecond scale, showcasing its capability for precise measurements.
@@
Figure 8: This represents the Schematic diagram for the MALTA testing layout illustrating the working flow of the experimental set-up. Upon instructions, FPGA triggers the laser and sends the actual triggering time to our oscilloscope. The laser falls on the X-ray tube, resulting in X-ray emission. The sensor detects the X-ray exposure and sends the detection time to the oscilloscope

### v2

@@ Abstract
The MALTA monolithic active pixel detector is developed to address some of the challenges anticipated in future high-energy physics detectors. As part of its characterization, we conducted timing studies necessary to provide a figure of merit for this family of monolithic pixel detectors. MALTA has a metal layer in front-end electronics, and the conventional laser technique is not suitable for timing studies due to the reflection of the laser from the metallic surface. X-rays have been employed as a more effective alternative for penetration through these layers. The triggered X-ray set-up is designed to study timing measurements of monolithic detectors. The timing response of the X-ray set-up is characterized using an LGAD. The timing response of the MALTA and MALTA2 pixel detectors is studied, and the best response time of MALTA2 pixel detectors is measured at about 2.6 ns.
@@ 1 Introduction
Pixel detectors are crucial in particle physics experiments due to their ability to precisely capture and record the individual locations of particle interactions [1]. Their high spatial resolution and fine granularity contribute to improved tracking, aid in analyzing complex event topologies, and enhance the overall performance of experiments in unraveling fundamental aspects of particle behavior and interactions. The increasing energy and luminosity of the Large Hadron Collider at CERN [2, 3], along with future experiments, require pixel detectors that can operate reliably in challenging radiation environments while maintaining high detection efficiency, low noise, minimal dead time, and low latency. Therefore, there is a constant effort to design efficient and thinner pixel detectors suitable for high-luminosity collisions. Monolithic silicon pixel detectors [5, 6, 7, 8, 9] offer several advantages over traditional bump-based hybrid detectors, including higher precision, lower noise, reduced material usage, and the potential for large-area production at lower cost. Despite their comparatively lower radiation hardness, these benefits have motivated significant R&D in CMOS-based pixel technologies. Such efforts have led to the development of the MALTA [10, 11] and its second-generation MALTA2 [12, 13] pixel detectors, specifically targeting applications in the HL-LHC upgrades and other future experiments. One important property of pixel detectors is fast timing, defined as the delay or the time taken between the interaction of a particle with the detector and the corresponding detection signal. In this study, a micro X-ray source is designed and employed to investigate the fast-timing performance of the MALTA and MALTA2 detectors.
2 MALTA and MALTA2 Demonstrator Systems
MALTA, designed using Tower-Jazz 180 nm technology [10], is a robust monolithic active pixel detector resilient to radiation exposure up to a fluence on the order of $10^{15}\,\text{n/cm}^{2}$ and a hit rate capability of up to $100\,\text{MHz/cm}^{2}$. Featuring an active area of $18.3\,\text{mm}^{2}$ and a $512\times 512$ pixel matrix with $36.4\times 36.4\,\mu\text{m}^{2}$ pixels, it boasts a quick pixel response time of up to $20\,\text{ns}$, meeting the HL-LHC’s bunch-crossing identification requirements. The chip’s asynchronous read-out eliminates the need for clock distribution, reducing power consumption. MALTA features a double-column read-out architecture and can read out multiple hits per double column. The chip is read out using a Virtex-7 series FPGA. Connection to a VC707 evaluation board is established through an FMC connector as shown in Fig 1(b), with a specially designed asynchronous oversampling detecting bits from the differential LVDS signals at an effective sampling rate of 1 Gsps, and subsequently read into the control software via IPBUS [14].
@@ 2 MALTA and MALTA2 Demonstrator Systems
MALTA2 [15, 16], a pixel detector demonstrator, is designed using Tower-Jazz 180 nm monolithic CMOS technology, measuring $20.2\,\text{mm}\times 10.1\,\text{mm}$. It is approximately half the size of MALTA, featuring a matrix of $224\times 512$ pixels, with each pixel measuring $36.4\times 36.4\,\mu\text{m}^{2}$. Two process modifications (NGAP and extra deep PWELL/XDPW) have been implemented to increase the charge collection efficiency in the pixel corners [17]. It is specifically designed to improve radiation hardness [18] while achieving lower random telegraph signal (RTS) noise in the sensor front-end. Improvements in MALTA2 slow control include using a shift register instead of an Ethernet-like protocol. In the pixel front-end, the implementation of open-loop amplification led to a compact design with low noise, and the use of multiple cascode transistors helped enhance the gain of the front end [12]. The demonstrator’s asynchronous read-out is clock-independent, transmitted through a 37-bit bus, and controlled via a 4322-bit shift register. The read-out uses a KC705, a KINTEX-7 series FPGA [19].
@@ 3 Micro-X-ray set-up at CERN
With the bunch-crossing clock of the LHC being 25 ns [20], sensors and read-out architecture have to be sufficiently fast in order to match the hits with the corresponding bunch-crossing. Furthermore, nanosecond-level response time resolution is required to contribute to event reconstruction targeting future HL-LHC upgrades. The fast timing of the chip requires examination prior to its usage in a detector. Traditionally, this investigation involves reflecting a laser beam off the detector and analyzing the discrepancy between the input and output signals [21, 22]. However, this laser technique proves inadequate for the MALTA2 due to significant reflection issues with the detector’s front-end electronics [10], which interfere with accurate timing measurements.
Figure 1: Figure illustrating the devices used for the study and the connections between them. In (a), the upper-left plot shows the complete set-up with a representation of a laser, a high-voltage source for X-ray generation, and an oscilloscope. In the left plot of (b), the Kintex-7 KC705 FPGA is shown, which is used to trigger the laser and also assists in signal processing. The right figure in (b) shows the MALTA sensor with the designed micro X-ray tube, held by a linear stage 10 cm above the chip.
To circumvent these challenges, we have opted for X-rays due to their ability to penetrate the metal layers of the detector without reflection. This does not require a high-end X-ray machine; instead, we have developed a custom-made micro X-ray system, as shown in Figure 1, which is capable of emitting X-rays in response to laser exposure. This approach employs a micro X-ray source that generates short pulses from Cu-Cr targets specifically designed for this application.
For the purpose of characterizing the timing performance of MALTA2, our triggered X-ray set-up plays an important role. Initially, the timing response of the X-ray itself is measured using an LGAD [23]. The LGAD used was a CNM LGAD with a pitch of $1.3\times 1.3\,\text{mm}^{2}$ and a gain of about 10–20, depending on the bias voltage applied [24]. The readout was done with discrete electronics, including a first and second stage amplifiers [24]. This baseline measurement serves as a reference for further comparison with MALTA2’s timing response, enabling us to achieve a comprehensive understanding of the detector’s timing capabilities.
@@ 3.1 Micro-X-ray Source
X-rays are generated by utilizing a 16 mm elongated tube featuring a photo-cathode and an anode composed of chromium and copper, fabricated above a Beryllium window. Electrons are emitted from the photo-cathode via the photoelectric effect initiated by a laser beam. Subsequently, these emitted electrons undergo acceleration through the application of a 12 kV electric field within the tube. The interaction of these accelerated electrons with the target material in the anode produces X-rays, including bremsstrahlung and fluorescence. The X-ray production process is illustrated in Figure 2. The emitted X-rays exhibit a random trajectory in space. With an increased applied voltage to the X-ray tube, the number of generated photons increases due to the formation of new transition lines.
Figure 2: Illustrative description of Micro X-ray source. It is connected to a 530 nm laser that falls on the cathode. The anode material is made with 250 nm copper and 50 nm chromium, which generates X-rays after interacting with accelerated electrons. Finally, there is a beryllium window that prevents electrons from passing through it.
@@ 3.2 Geant4 Simulation for X-ray tube
Figure 3: Geant4 simulation of photon energies generated from the micro X-ray tube. The x-axis represents the photon energy in keV. Four characteristic peaks are visible: two prominent peaks at 5.41 and 8.04 keV corresponding to the $K_{\alpha}$ lines of Cr and Cu, and two smaller peaks at 5.95 and 8.90 keV corresponding to their $K_{\beta}$ lines, respectively.
A Geant4 simulation was conducted to analyze the energy variation of photons emitted from an X-ray tube. The simulation results in Figure 3 show distinct peaks at 5.41 and 8.04 keV, corresponding to the characteristic $K_{\alpha}$ lines of chromium and copper, along with two distinct $K_{\beta}$ lines. Subsequently, experimental validation of these findings was undertaken using an LGAD (Low Gain Avalanche Detector) because of its high granularity and precise timing capabilities.
@@ 4 Test Set-up with LGAD
The LGAD, or Low-Gain Avalanche Detector [23, 24, 25], is a type of silicon detector used in particle physics experiments. The utilization of LGAD in this experimental phase is particularly advantageous, as it enables optimal timing resolutions [26], a crucial factor in accurately characterizing X-ray emission.
The LGAD experimental arrangement is illustrated in the accompanying schematic diagram, shown in Figure 4. In this set-up, laser beams are generated and guided onto a micro X-ray source, initiating the generation of X-rays. The resultant X-rays are directed toward the LGAD, and the corresponding signals are captured and recorded using an oscilloscope. Simultaneously, the LGAD signal is directly used to access the energy of the emitted photons based on the amplitude of the LGAD signal. This approach provides a direct and efficient means of evaluating the photon energy within our experimental framework.
@@
Figure 4: LGAD set-up illustrating the working flow of the experiment. In (a), the laser is focused directly on the LGAD, and in (b), an additional micro X-ray source is used. The laser emission and detection time information are collected and compared in both cases using an oscilloscope to extract individual contributions from the laser and the X-ray source.
@@ 5 Results from LGAD
We conducted two key studies utilizing the Low-Gain Avalanche Diode. The first study focused on characterizing the energy peaks generated by the X-ray source, which is necessary to understand the X-ray tube’s performance and ensure its suitability for our applications. The second study analyzed the timing precision of the X-ray source by examining the standard deviation of its delay. For this purpose, we compared the performance of the micro X-ray source, activated by laser exposure, against a standalone laser source. These preliminary analyses are important for evaluating the contribution of the X-ray set-up to the total delay, ensuring precise timing measurements in our experimental framework.
@@ 5.1 Characterization of Micro X-ray Tube
Figure 5: Comparison of signal amplitudes measured using the LGAD with an $^{55}\mathrm{Fe}$ source and the X-ray tube as a function of voltage applied to the LGAD. Two peaks are visible for the signal using X-rays generated at a 12 kV supply voltage, corresponding to the Cr and Cu $K_{\alpha}$ lines. The peaks of the red line, representing the amplitude of the signal using the X-ray tube at 12 kV, are compared with the peak from the $^{55}\mathrm{Fe}$ source (blue).
In this study, we quantify the energy of photons using the LGAD signal, represented on the x-axis in terms of voltage and exhibiting a linear relationship with energy. For calibration, we employed a $^{55}\mathrm{Fe}$ source, which emits $K_{\alpha}$ X-rays with a well-known energy of 5.9 keV, thereby establishing the calibration framework. The measured energy from the alpha decay was subsequently compared with the energies of X-rays generated at voltages of 5.5 kV,  8.5 kV,  and 12 kV.  Notably, the X-rays produced under a 12 kV voltage setting revealed distinct peaks at 5.4 keV and 8 keV, which correspond closely to the theoretical $K_{\alpha}$ emission lines of chromium and copper, respectively. These peaks occur because the incident electron energy at 12 kV exceeds the K-shell binding energies of these elements, allowing characteristic X-ray lines to appear alongside the bremsstrahlung continuum. As the energy of the X-ray photon increases, the emitted electron gains more energy, resulting in a higher signal voltage. A comparison of peaks is shown in Figure 5. The peak is parametrized using Gaussian and polynomial functions, as shown in Figure 6. The results from the fit are summarized in Table 1, considering the proportional values of output signal voltages with X-ray energy, which match the theoretical values of $K_{\alpha}$ lines from chromium and copper.
Figure 6: Comparison of the fitting parameters of generated signal amplitudes using the LGAD with $^{55}\mathrm{Fe}$ and X-ray sources at 5.5 kV, 8.5 kV, and 12 kV. The x-axis represents the energy of generated electrons from the LGAD using the sources, while the y-axis represents the number of events. Two peaks are visible for the signal amplitude using X-rays generated at 12 kV supply voltage. These plots are fitted using Gaussian(s) for peaks and a polynomial function to fit the initial points.
@@
Peak
Value
(in V)
Energy
(Theoretical value)
Energy
(Observed value)
@@
Peak from 12 keV Cr 0.0188 5.41 keV 5.54 keV
Peak from 12 keV Cu 0.0270 8.04 keV 7.96 keV
@@ 5.2 Timing performance of LGAD
In this investigation, the LGAD was characterized using a laser beam, yielding timing resolutions in the range of 25–50 ps, where the resolution includes contributions from both the LGAD and the laser source. Subsequently, the entire system was employed to generate X-rays through laser irradiation. From the analysis of the X-ray signals, the timing resolution was determined to be approximately 270–300 ps across varying laser frequencies, as shown in Figure 7. This evaluation allowed us to infer that the intrinsic timing resolution of the micro X-ray source is estimated to be in the range of 269–295 ps ($\sqrt{270^{2}-25^{2}}$$\sqrt{300^{2}-50^{2}}$ ps), demonstrating its efficiency and reliability in temporal measurements. These findings highlight the potential of our system for applications requiring precise timing resolution in X-ray generation and detection on the order of 100 ps.
Figure 7: Comparison of the timing delay between the laser trigger and LGAD signal detection using the laser and X-rays. The laser frequency is varied from 5 to 15 MHz, and the RMS of the delay is compared without (Figure 4(a)) and with (Figure 4(b)) the micro X-ray source. The RMS of the timing delay using the micro X-ray tube with the laser is found to be approximately 250 ps higher than the RMS of the timing delay without the micro X-ray tube, indicating the contribution of the micro X-ray tube to the RMS of the delay.
@@ 6.1 Measurements Using MALTA Sample
In contrast to the LGAD set-up, the MALTA system employs an internal trigger from the FPGA to initiate the laser, which is subsequently directed toward the X-ray tube for X-ray emission, as described in Figure 8. The operational sequence involves triggering the FPGA using software, whereby the KINTEX-7 FPGA activates the laser source. As the laser beam interacts with the X-ray tube, X-rays are emitted and directed toward the MALTA sensor (W11R12, a 100 nm thick sensor with a p-type Czochralski substrate and an extra deep p-well implant). The timing information is gathered within the FPGA and transmitted to the oscilloscope, which then compares the timing data to determine the delay. This delay information is necessary for analyzing the temporal characteristics of MALTA, where the effect of the previously measured micro X-ray source (approximately 300 ps) can be adjusted. Notably, the timing performance of MALTA is observed to operate on a nanosecond scale, showcasing its capability for precise measurements.
@@
Figure 8: Schematic diagram of the MALTA testing layout illustrating the working flow of the experimental set-up. Upon instruction, the FPGA triggers the laser and sends the actual triggering time to the oscilloscope. The laser falls on the X-ray tube, resulting in X-ray emission. The sensor detects the X-ray exposure and sends the detection time to the oscilloscope.